OpenStudy (anonymous):

Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. x > –2 y ≤ 2x + 7 HELP ME PLEASE WILL MEDAL AND FAN

OpenStudy (anonymous):

It helps to pretend these are equations (not inequalities) at first. Graph x = -2 (a vertical line at the point x = -2). However, because it's > and not "> or =", it should be a dotted line, not a solid line. Now...where are things "Greater than" x = -2? That's where you need to shade... x > -2 is everything to the right of your line, so shade that area. Now, graph y = 2x + 7. (Note, because this is "< or =" it will be a solid line.) There are a few ways to graph this line - all of which should be familiar to you by now. The easiest way is to put a dot on the y-intercept (0, 7), then find the next point by using the slope of 2. So, another point would be (1, 9). To find the area that is "< or =" to the line y = 2x + 7, you need to pick a point that is NOT on that line and see if that point is true or false. I'll pick (1, 1). y <= 2x + 7 1 <= 2(1) + 7 1 <= 9 That is true, so shade the graph on the side of the line that contains that point. (If it was FALSE, you would shade on the side of the line that does NOT contain that point.) The area where the two shadings overlap is the solution to the system of inequalities. (It's more than one point.) The point where the two lines intersect (-2, 3) is the solution to the system of equations (x = -2 and y=2x+7), not inequalities.

OpenStudy (anonymous):

The answers to choose from are (–3, –1) (–1, 4) (1, 11) (–1, 6) @MW3030

OpenStudy (anonymous):

@imqwerty @pyxieyt

OpenStudy (anonymous):

sorry i cant figure it out and i dont want to give you the wrong answer

OpenStudy (anonymous):

@MW3030 that's okay! thank you for trying.

OpenStudy (anonymous):

@pyxieyt @imqwerty HELP

OpenStudy (pyxieyt):

It helps to pretend these are equations (not inequalities) at first, okay?

OpenStudy (pyxieyt):

Graph x = -2 (a vertical line at the point x = -2). However, because it's > and not "> or =", it should be a dotted line, not a solid line.

OpenStudy (anonymous):

okay.

OpenStudy (pyxieyt):

Now...where are things "Greater than" x = -2? That's where you need to shade... x > -2 is everything to the right of your line, so shade that area.

OpenStudy (pyxieyt):

Now, graph y = 2x + 7. (Note, because this is "< or =" it will be a solid line.) There are a few ways to graph this line - all of which should be familiar to you by now.

OpenStudy (anonymous):

okay I see.

OpenStudy (pyxieyt):

The easiest way is to put a dot on the y-intercept (0, 7), then find the next point by using the slope of 2. So, another point would be (1, 9).

OpenStudy (anonymous):

okay.. Um the closest I get is 1,6

OpenStudy (pyxieyt):

To find the area that is "< or =" to the line y = 2x + 7, you need to pick a point that is NOT on that line and see if that point is true or false.

OpenStudy (anonymous):

wonder how you all are getting that. I already posted the answer choices they are nothing like that.

OpenStudy (pyxieyt):

His steps are correct, that's why I'm typing it out. He just used that as an example

OpenStudy (pyxieyt):

You could do what he did with the example to find out.

OpenStudy (pyxieyt):

(I just found what he was looking at,it was yahoo :( )

OpenStudy (anonymous):

okay.. I looked up the question on yahoo answers.com and it shows the exact same thing.. But the answers are totally different.

OpenStudy (pyxieyt):

Yeah so I think he looked it up.

OpenStudy (pyxieyt):

Let's plug in the answers that are above.

OpenStudy (anonymous):

okay let me know what u get. Algebra is har.

OpenStudy (pyxieyt):

I'll pick (–3, –1) y <= 2x + 7 -1 <= 2(-3) + 7 1 <= 1 Do you think this is correct?

OpenStudy (anonymous):

probably. I will put it as the answer and see what I get.

OpenStudy (pyxieyt):

Wait! we need to go through it all

OpenStudy (anonymous):

ok

OpenStudy (pyxieyt):

Remember, because you may think it is one answer but it may not be!

OpenStudy (anonymous):

true so do the rest. :)

OpenStudy (pyxieyt):

Can you try to put in the next one I'll pick (–1, 4). y <= 2x + 7 ? <= 2(?) + 7 1 <= ?

OpenStudy (pyxieyt):

Remember I still need you to try and cooperate. If you told try it, it would be like i'm just giving you the answers!

OpenStudy (pyxieyt):

*dont not told xD

OpenStudy (anonymous):

Yes I know. Do I plug in every number for x until I get the answer.

OpenStudy (pyxieyt):

Yes put the x coordinates in the x variable and the Y coordinates in the Y variable to see if the equation is true.

OpenStudy (pyxieyt):

I will check your work to see if it is correct.

OpenStudy (anonymous):

okay one minute, I am looking to see if I can find it somewhere else

OpenStudy (pyxieyt):

No you should figure it out yourself, if you don't how would you learn?

OpenStudy (anonymous):

I figured it out right before I clicked a link! It's (-1,4)

OpenStudy (pyxieyt):

Lets try B. together! -1 > –2 I'll pick (–1, 4). y <= 2x + 7 4<= 2(-1) + 7 4 <= 5 Yep!

OpenStudy (anonymous):

So, B or my answer?

OpenStudy (pyxieyt):

Yes B is your answer! Remember next time try not to look it up on yahoo! On OpenStudy we can help guide you through it and help you learn!

OpenStudy (anonymous):

okay thank you @pyxieyt